OpenStudy (anonymous):

Given f(x)= ln(x-1) a. find the equation (in the y=mx+b form) of the tangent to f(x) at x=3 b. find the zeros of f(x) analytically c. sketch the graphs of f(x) and the tangent line on the same axes d. write your procedure for finding the zeroes of f(x) graphically PLEASE HELP AND SHOW ALL STEPS!

OpenStudy (anonymous):

a)f(x)= ln(x-1) f '(x)= 1/( x-1) f '(3)= 1/(3-1)=1/2, also f(3)= ln(3-1)=ln 2, using y-y1=m(x-x1) y-ln 2=(1/2)(x-3), therefore y=0.5(x-3)+ln2 b)f(x)= ln(x-1)=0 e^ln(x-1)=e^0 X-1=1 X=2 ans....